# Assurance of AI Systems From a Dependability Perspective

### v2

@@ 3 Assurance of AI Systems for Specific Functions
Another architecture for diverse redundancy is the “dual-process” approach similar to that described in the previous section, where the output of an ML process (typically, one performing perception) is presented to a symbolic AI process supplied with some generic (or “common sense”) model of the environment. This may be able to reject some erroneous perceptions and draw deeper conclusions from plausible ones [136]. For example, a car’s perception system operating in a “freeway ODD” that detects a bicycle may be overridden by a second-level symbolic AI that “knows” bicycles are not allowed on freeways. Although experimental or theoretical confirmation are lacking, it does seem plausible that the dual processes would be diverse and might be expected to fail independently.
@@
A particular attraction of such dual process architectures is that the second (upper) process does have a (generic) model and compensates for the model-free behavior of the first. A more sophisticated dual-process architecture, as sketched in the previous section and inspired by that of the human brain, uses predictive processing. Here, the two processes combine to build an explicit model of the world at runtime [^22] We should explain how a model-free perception system is used to build a world model. The object detector of a self-driving car, say, has no model of what it is seeing: it does not “know” that it is on a road and looking at other traffic; it has simply been trained to recognize objects (albeit using billions of road images). Hence it may report a “boat” when “trailer” (with a boat as cargo) would be more useful. The upper levels of the perception pipeline are explicitly designed to build a model of traffic on a road, and apply this built-in interpretation to the model-free output of the object detector. and, rather then rely directly on the model-free products of the lower process, this highly specific model is used to predict them and they are therefore interpreted in this light; the higher-level process intervenes when a large prediction error indicates “surprise.”
@@
In some circumstances we can use human reviewers or supervisors rather than automated processes to provide assurance: many AI-based systems work collaboratively with, or under the supervision of, human operators who might be expected to provide trustworthy runtime checks. However, there are well-known concerns (the “ironies of automation” [12]) about human attention and responsibility in these circumstances: typically, the human either ignores the automation or trusts it completely (a phenomenon known as automation bias) [158, 52]. Moreover, artificial “intelligence” may be too narrow to engage effective human collaboration and interaction. For although AI may outperform humans on some tasks, human cognition combines and integrates many capabilities beyond those of current AI, including realtime learning and memory, reasoning, analogy, abstraction, generalization, and planning. In addition, these capabilities employ and integrate a diverse range of sensory inputs, including sound, smell, touch, vibration, and many others, that provide wide situation awareness. AI systems can have limited approximations to some of these capabilities (e.g., elementary planners; runtime ML for learning and memory with, possibly, some generalization [181]; and automated deduction for reasoning) but these are weakly integrated and typically rely on specifically focused sensors. Furthermore, AI generally lacks higher-order capabilities (i.e., thinking about thinking, sometimes referred to as metacognition) that are needed for longer term planning and for situation awareness in uncertain environments.
@@
As a result of their different capabilities, humans and AI will build different models of the world and will generate and interpret their goals and inputs differently (and, as we have seen, those based on LLMs may have a model-free foundation). Consequently, humans may be poor judges of automation behavior and vice versa. Individual humans also build different models so, when they communicate, they compensate by using a “theory of mind,” which is a model of the other participants’ state of knowledge, beliefs, desires, intentions, and so on [5]. To be assurable or checkable by humans, AI must do something similar: it is not enough to employ “Explainable AI” to describe the internal details of its own “reasoning,” it must do so relative to an accurate theory of mind for its interlocutors [162]. The theory may need to hypothesize that the other party has an incorrect model or false beliefs. For example, there are numerous airplane crashes caused by pilots misinterpreting their situation, usually by fixating on one indicator, and doing the wrong thing (e.g., shutting down the good engine). [^23] The Kegworth crash of British Midland Flight 92 on 8 January 1989 is instructive. The left engine suffered a broken fan blade, filling the cabin of the 737-400 with smoke. The pilots were familiar with earlier models of the 737 where the cabin ventilation “packs” are on the right engine (whereas on the 400 they are on the left) so they assumed the faulty engine was on the right (because of the smoke) and shut it down while increasing power on the left engine, which then failed, leading to the crash and loss of 47 lives. Several instruments correctly indicated the left engine as the faulty one, as did the passengers and flight attendants, but the pilots persisted with their flawed model. Thus, an AI “co-pilot” must diagnose not only the physical fault but also the pilot’s faulty mental model and should then attempt to direct their attention to the contrary indicators [3, Appendix A].
@@
In summary, we have seen that assurance for AI systems designed for specific functions begins by identifying hazards and then seeking ways to mitigate those hazards. Sometimes this can be achieved by selection of ML training data, or by monitoring explanations, or by comparing behavior for variant inputs. In some circumstances, it may be feasible to construct verifiable guards, but often the guards must use AI and ML in their own perception systems, thereby vitiating strong assurance. Even so, it may be feasible to achieve modest levels of assurance based on diversity, providing the perception system (or whatever builds the system’s world model) also has a diverse architecture. Another possibility is human supervision, but this raises issues of mutual understanding and other challenging topics in human-computer interaction.
A legitimate question is how can we build indefeasible assurance cases for systems with AI and ML where the mechanisms and evidence for assurance are relatively weak? First, it is important to remember that indefeasible does not mean inerrant (i.e., “no failures”), it means no new information would change our judgement, and this in turn means that we have diligently searched for all hazards and all defeaters to the assurance case, and have dealt with them. But in systems with AI and ML, the total space of behaviors is unknown and so the true failure rate cannot be determined, The best that we can do is to show that the failure rate is above some conservative “floor” and that if experience reveals previously unanticipated misbehavior, then we will detect and correct it before it causes unacceptable harm.
Hence, the top claim of a system with AI and ML elements will not be “no catastrophic failure in entire deployment lifetime” (as it is for aircraft software) but something like “incident rate caused by the system shall be no more than $10^{-n}$ per hour, and potential violations experienced in operation will be detected and mitigated before a second incident.”
Such a claim will be established by evidence that includes analysis and experimentally-based reliability estimation to justify the modest “floor” plus robust fleet-wide methods to detect and record incipient failures during operation (e.g., following a “surprising” prediction error or, in the worst case, an accident) and to rapidly develop and distribute repairs (recall the discussion in Section 1.2 of repairs following incidents and precursors). Mechanisms for runtime verification such as guards are the natural location for assessing dependability weakness and detecting incipient failures, which Johansson and Koopman refer to as Safety Performance Indicators (SPI) [86]. Notice that these guards are not only monitoring the system, but also collecting evidence that will update the assurance case; we refer to this architecture as providing a dynamic assurance case. Note that it may be reasonable to use strict dependability assurance for some properties (e.g., forward collisions) and dynamic assurance for others (e.g., sideways collisions).
@@ 4 Assurance for General-Purpose AI
By general-purpose AI we mean those AI and ML systems that are broadly capable and can be adapted to some specific application with relatively little effort. Canonically, these are generative “foundation models” such as LLMs (for language) or “generative adversarial networks” (GANs) and “diffusion models” (for images) that are pre-trained using unsupervised learning on vast amounts of unlabeled data so that they learn general features and associations rather than specific skills. They can be adapted to specific tasks using “prompts,” and “prompt engineering” has become a recognized activity [127]. We refer to such systems generically as LLMs and we discussed assurance for applications built on them in the previous section; here, we focus on the LLMs themselves.
@@
Unlike applications for specific purposes where explicit hazards can be identified and mitigated, general-purpose AI and ML tools present more of a challenge to assurance because the hazards will depend on how they are employed, so any protections need to be similarly general-purpose and to apply in all circumstances. Furthermore, since a general-purpose tool will support many applications, it will be subjected to more demands than any single application and should therefore be more highly assured, even though its assurance seems more difficult.
@@
One manifestation of this generality and difficulty is that public discourse and much technical literature speaks of trust rather than assurance. Indeed, many current AI systems are sufficiently good at natural language, including emotional and cultural nuances, that people often anthropomorphize them [^24] For example, in June of 2022, Blake Lemoine, a Google engineer working with their large language model, LaMDA (Language Model for Dialogue Applications), claimed it had become “sentient” (Washington Post, 11 June 2022). and spontaneously endow them with trust. In fact, the developers of general-purpose LLMs engineer them to encourage, and to some extent earn, this trust, as we now outline.
@@
A bare LLM, trained to predict the next or missing “token” in a string of text, is as likely to produce wrong or noxious results as good ones. This is because its training uses a massive and largely indiscriminate sampling of material found on the Web, much of which is false or noxious. The bare LLM is then “protected” (from generating noxious or potentially harmful output) using reinforcement learning where human reviewers provide feedback on candidate results [11]. This technique has been developed into a more automatic process by first providing the LLM with a list of principles, referred to as its “constitution,” and then using reinforcement learning with the feedback on alignment between its responses and its constitution provided by a second AI agent (a process referred to as “Reinforcement Learning with AI Feedback”, RLAIF) [10]. The constitution used with Anthropic’s LLM Claude is available online https://www.anthropic.com/news/claudes-constitution. It consists of a few dozen prompts, including the following examples.
@@
Because LLMs typically present humanoid “chatbot” personas, public—and even some technical—assessments of their trustworthiness are of the interpersonal variety and resemble those applied to humans [107]. However, this trust does not rest on assured technical grounds and we cannot predict when its mechanisms will be effective and when they will fail. For example, the protective mechanisms of Google Gemini’s image generation distorted its behavior and exposed it to widespread ridicule [135]. And studies by the UK’s AI Safety Institute found that safeguards on LLMs are largely ineffective. [^26] See https://www.aisi.gov.uk/work/advanced-ai-evaluations-may-update. Also note that recent demonstrations show how LLMs can be crafted to subvert constraints imposed upon them [78] and that fine-tuning for specific purposes can unwittingly compromise protections [134].
@@
We will interpret the rules and constraints underlying social trust as contracts and thereby equate social trust with contractual trust. If we further interpret failure to uphold a contract as a hazard, and “vulnerability” as indicating that such failures impose costs on the trustor, then contractual trustworthiness looks very much like dependability assurance with the contract as its requirements. For this reason, we regard contractual trust in AI as an issue of dependability assurance. However, contractual trust requires a contract and these are more readily understood with the focused AI systems of Sections 2 and 3 (where the contract corresponds to their dependability requirements) than with the general-purpose systems considered here.
In fact, there are two issues here: a) what constitutes a generically useful contract, and b) how it is enforced. Generic contracts will be less focused on hazards, since those are specific to applications, and more on general concerns including alignment with human values such as fairness and honesty. Claude’s “constitution” is one model for a generic contract, but for assurance we would want it to be enforced explicitly by an assured external guard rather than implicitly by the LLM itself. Hence, we will next examine very general contracts or constraints for AI systems, and methods for enforcing them. As frameworks, we will consider ways that human behavior is constrained: that is, by ethical and legal standards, supervision, rewards and punishment, reputation, and so on. Some of this material is drawn from a previous report on controls for potentially conscious agents [150].
@@
Ethical and related principles are referred to as normative, and symbolic AI systems have been developed that can represent such principles and thereby perform “normative reasoning” [41]. These have been applied to trolley problems, including some that involve self-harm (e.g., throwing yourself in front of the trolley) and thereby violate the “unintended” aspect of DDE [32, 65]. It is claimed that fairly sophisticated logical treatments (e.g., intensional logics, counterfactuals, deontic modalities) are needed to represent normative scenarios, and these might be additional to what is needed for the primary functions of the system (hence, must be introduced explicitly, which will add complexity). Additionally, when normative requirements are introduced as rules in some formal notation, there is concern that they may have internal conflicts or otherwise lack wellformedness, and methods have been proposed for checking this [58].
Other recent work formalizes Kant’s categorical imperative (humans must be treated as ends, not as means), which requires a treatment of causality [110], while others favor “Virtue Ethics” derived from Aristotle [170]. And note that we have mentioned only logical or rule-based representations for ethics, whereas game theory provides another perspective.
Application of constraints derived from general normative principles requires the guard to build a model of its world and to interpret the principles appropriately for that world. It is unlikely that a non-AI perception system can build a world model that is adequate for this task, so the world model must be built by an AI perception system, or even based on that claimed by the LLM itself. We have seen previously that assurance for world models constructed by AI-based perception systems is challenging and currently infeasible, so only weak forms of safety assurance can be delivered by guards based on ethics and other general rules.
@@
Building an accurate and comprehensive model of the world is difficult in a focused application; it is next to impossible for a general-purpose tool lacking information on the possible contexts of its use.
@@
We conclude that general constraints built into guards for general-purpose AI tools cannot provide automated protection nor assurance, largely because their correct application depends on the guard building an accurate model of the world in which to interpret them, or on trusting the LLM itself to provide a suitable model (which is contrary to its generally model-free operation). It is, however, possible that, with suitable programming interfaces or suitable sensitivity to prompts, the presence of general constraints could provide useful capabilities that humans can use to add controls for specific applications built on these general-purpose tools, albeit with weak assurance.

### v3

@@ 3.4 Diversity
Another architecture for diverse redundancy is the “dual-process” approach similar to that described in the previous section, where the output of an ML process (typically, one performing perception) is presented to a symbolic AI process supplied with a domain model of the environment, plus a generic one providing “common sense.” This may be able to reject some erroneous perceptions and draw deeper conclusions from plausible ones [193]. For example, a car’s perception system operating in a “freeway ODD” that reports a bicycle (perhaps painted on the back of a truck) may be overridden by a second-level symbolic AI that “knows” bicycles are not allowed on freeways. Although experimental or theoretical confirmation are lacking, it does seem plausible that the dual processes would be diverse and might be expected to fail independently.
@@
A particular attraction in using such dual process architectures with LLMs is that the second (upper) process does have a domain model and compensates for the model-free behavior of the lower-level LLM. A more sophisticated dual-process architecture, as sketched in the previous section and inspired by that of the human brain, uses predictive processing. Here, the two processes combine to build an explicit local model of the world and, rather then base it directly on the model-free outputs of the lower process, the local model is used to predict those outputs and they are interpreted in this light; the domain model of the higher-level process intervenes when a large prediction error indicates “surprise” [42]. LeCun’s Joint Embedding Predictive Architecture (JEPA) shares some of these characteristics [152, 86].
3.5 Human Review
@@
In some circumstances we can use human reviewers or supervisors rather than automated processes to provide assurance: many AI-based systems work collaboratively with, or under the supervision of, human operators who might be expected to provide trustworthy runtime checks. However, there are well-known concerns (the “ironies of automation” [14]) about human attention and responsibility in these circumstances: typically, the human either ignores the automation or trusts it completely, a phenomenon known as automation bias [221, 73]. Moreover, artificial “intelligence” may be too narrow to engage effective human collaboration and interaction. For although AI may outperform humans on some tasks, human cognition combines and integrates many capabilities beyond those of current AI, including realtime learning and memory, reasoning, analogy, abstraction, generalization, and planning. In addition, these capabilities employ and integrate a diverse range of sensory inputs, including sound, smell, touch, vibration, and many others, that provide wide situation awareness. AI systems can have limited approximations to some of these capabilities (e.g., elementary planners; runtime ML for learning and memory with, possibly, some generalization [250]; and chain of thought plus automated deduction for reasoning) but these are weakly integrated and typically rely on specifically focused sensors. Furthermore, AI generally lacks higher-order capabilities (i.e., thinking about thinking, sometimes referred to as metacognition) that are needed for longer term planning and for situation awareness in uncertain environments.
@@
As a result of their different capabilities, humans and AI will build different models of the world and will generate and interpret their goals and inputs differently (and, as we have seen, those based on LLMs may have a model-free foundation). Consequently, humans may be poor judges of automation behavior and vice versa. Of course, individual humans also build different models so, when they communicate, they compensate by using a “theory of mind,” which is a model of the other participants’ state of knowledge, beliefs, desires, intentions, and so on [6]. To be assurable or checkable by humans, AI must do something similar: it is not enough to employ “Explainable AI” to describe the internal details of its own operation, it must do so relative to an accurate theory of mind for (i.e., from the point of view of) its interlocutors [225]. The theory may need to hypothesize that the other party has an incorrect model or false beliefs. For example, there are numerous airplane crashes caused by pilots misinterpreting their situation, usually by fixating on one indicator, and doing the wrong thing (e.g., shutting down the good engine). [^23] The Kegworth crash of British Midland Flight 92 on 8 January 1989 is instructive. The left engine suffered a broken fan blade, filling the cabin of the 737-400 with smoke. The pilots were familiar with earlier models of the 737 where the cabin ventilation “packs” are on the right engine (whereas on the 400 they are on the left) so they assumed the faulty engine was on the right (because of the smoke) and shut it down while increasing power on the left engine, which then failed, leading to the crash and loss of 47 lives. Several instruments correctly indicated the left engine as the faulty one, as did the passengers and flight attendants, but the pilots persisted with their flawed model. Thus, an AI “co-pilot” must diagnose not only the physical fault but also the pilot’s faulty mental model and should then attempt to direct their attention to the contrary indicators [3, Appendix A].
3.6 Summary for AI in Systems for Specific Functions
We have seen that assurance for AI systems designed for specific functions begins by identifying hazards (which may be novel) and then seeking ways to mitigate those hazards.
@@ 3.6 Summary for AI in Systems for Specific Functions
Sometimes this can be achieved by selection of ML training data, or by monitoring explanations, or by comparing behavior for variant inputs. In some circumstances, it may be feasible to construct verifiable guards, but often the guards must use AI and ML in their own perception systems, thereby vitiating strong assurance. Even so, it may be feasible to achieve modest levels of assurance based on diversity, providing the perception system (or whatever builds the system’s world model) also has a diverse architecture. Another possibility is human supervision, but this raises issues of mutual understanding and other challenging topics in human-computer interaction.
We outline options for assured dependability in terms of the three classic strategies as follows.
Fault avoidance.
This cannot be a significant strategy for systems based on AI and ML as their development methods do not provide sufficient insight into their behavior, nor support for guarantees on their performance.
Fault tolerance.
This is a key strategy when components are relatively unreliable or hard to evaluate.
It is applicable to systems where AI has specific functions as these support hazard analysis and provide safety and security requirements.
The architecture patterns of the previous section are relevant and can support various forms of guards and checks, including those based on explanations, verification of outputs, dual-process monitoring with diversity, and human review.
Failure Management and Resilience.
As with conventional systems, resilience can be an effective strategy that can support “learning by doing” when the costs of failure are tolerable.
Traditionally, AI and ML systems were targeted at specific applications, such as radiology, and were trained on data for their particular application. Increasingly, however, general-purpose AI and ML foundation models are becoming broadly capable and are being applied, at least experimentally, in many different application domains. Accordingly, in the next section, we will examine assurance for general-purpose AI.
@@ 4 Assurance for General-Purpose AI
We now extend the discussion to general-purpose AI that can be deployed in a wide range of systems and contexts. By general-purpose AI we mean those AI and ML systems that are broadly capable and can be adapted to some specific application with relatively little effort. Canonically, these are generative “foundation models” such as LLMs (for language) or “generative adversarial networks” (GANs) and “diffusion models” (for images) that are pre-trained using unsupervised learning on vast amounts of unlabeled data so that they learn general features and associations rather than specific skills. They can be adapted to specific tasks using “prompts,” and “prompt engineering” has become a recognized activity [182]. We refer to such systems generically as LLMs (which are becoming “multi-modal,” meaning they are trained on images and sound as well as text) and we discussed assurance for applications built on them in the previous section; here, we focus on the LLMs themselves.
The characteristics of general-purpose AI bring some additional challenges.
  • Due to widespread applications, their impact is wide ranging with potential for unintended consequences.
  • Their facility with language can allow their system to become deeply and widely embedded in the social context, increasing the importance of sociotechnical concerns.
  • Previously, ethical issues would be addressed as part of the risk, hazard, and requirements analysis for each specific system, whereas general-purpose AI may need a more comprehensive treatment of ethics.
  • General-purpose AI must be assured independently of specific applications or contexts.
  • But there should also be a methodology with a supporting engineering process to adjust, constrain, and assure behavior when the specific application context becomes known.
We now elaborate on these issues, beginning with the first bullet.
We drew attention to system failures, Normal Accidents, and Self-Organizing Criticality in Section 1.1, and noted that these often arise from unanticipated interactions in complex systems. Systems built around general-purpose AI with interactions deeply embedded in their social context are candidates for this kind of complexity and vulnerability. However, as we noted in that earlier section, current general-purpose AI systems are sufficiently unreliable that they are a significant source of component failure in systems that employ them, and that will be our focus in this section.
We see two use cases for general-purpose AI: one is where the LLM is the system, or is thought of as such, as in chatbots where human users interact with the LLM, perhaps in a structured way, rather as if it were a colleague; the other is where the LLM is explicitly used as a component in a larger system. For the latter case, we discussed assurance for the overall system in the previous section, here we consider what should be done in and around the LLM to render it safer or more assurable in its role as a component. We suggest these considerations also apply to the chatbot case, but note that the true system boundary may be much larger than the chatbot itself and therefore system failures should be anticipated as well.
@@
Unlike applications for specific purposes where explicit hazards can be identified and mitigated, general-purpose AI and ML tools present more of a challenge to assurance because the hazards will depend on how they are employed, so any built-in protections need to be similarly general-purpose and to apply in all circumstances. Furthermore, since a general-purpose tool will support many applications, it will be subjected to more demands than any single application and should perhaps be more highly assured, even though its assurance seems more difficult.
@@
One manifestation of this generality and difficulty is that public discourse and much technical literature speaks of trust rather than assurance. Indeed, many current AI systems are sufficiently good at natural language, including emotional and cultural nuances, that people often anthropomorphize them [^24] For example, in June of 2022 a Google engineer working with their large language model, LaMDA (Language Model for Dialogue Applications), claimed it had become “sentient” (Washington Post, 11 June 2022). and spontaneously endow them with trust. In fact, the developers of general-purpose LLMs engineer them to encourage, and to some extent earn, this trust, as we now outline.
4.1 Trust
@@
A bare LLM, trained to predict the next or missing “token” in a string of text, is as likely to produce wrong or noxious results as good ones. This is because its training uses a massive and largely indiscriminate sampling of material found on the Web, much of which is false or noxious. The bare LLM is then “protected” (from generating noxious or potentially harmful output) using reinforcement learning where human reviewers provide feedback on candidate results [13]. This technique has been developed into a more automatic process by first providing the LLM with a list of principles, referred to as its “constitution,” and then using reinforcement learning with the feedback on alignment between its responses and its constitution provided by a second AI agent (a process referred to as “Reinforcement Learning with AI Feedback”, RLAIF) [12]; even more automatic methods have recently been developed [94]. The constitution used with Anthropic’s LLM Claude is available online at https://www.anthropic.com/news/claudes-constitution. It consists of a few dozen prompts, including the following examples.
@@
Because LLMs often present humanoid “chatbot” personas, public—and even some technical—assessments of their trustworthiness are of the interpersonal variety and resemble those applied to humans [156]. However, this trust does not rest on assured technical grounds and we cannot predict when its mechanisms will be effective and when they will fail. For example, the protective mechanisms of Google Gemini’s image generation distorted its behavior and exposed it to widespread ridicule [192]. And studies by the UK’s AI Safety Institute found that safeguards on LLMs are largely ineffective. [^26] See https://www.aisi.gov.uk/work/advanced-ai-evaluations-may-update. Also note that recent demonstrations show how LLMs can be crafted to subvert constraints imposed upon them [108] and that fine-tuning for specific purposes can unwittingly compromise protections [191].
@@ 4.2 Social Trust
We will interpret the rules and constraints underlying social trust as contracts and thereby equate social trust with contractual trust. If we further interpret failure to uphold a contract as a hazard, and “vulnerability” as indicating that such failures impose costs on the trustor, then contractual trustworthiness looks very much like assurance for dependability with the contract as its requirements. For this reason, we regard contractual trust in AI as an issue of dependability assurance. However, contractual trust requires contracts and these are more readily understood with the focused AI systems of Sections 2 and 3 (where contracts corresponds to their dependability requirements) and needs interpretation for the general-purpose systems considered here.
There are really two issues here: a) what constitutes a generically useful contract, and b) how it is enforced. Generic contracts will be less focused on hazards, since those are specific to applications, and more on general concerns, including alignment with human values such as fairness and honesty. Claude’s “constitution” is one model for a generic contract, but for assurance we would want it to be enforced explicitly by an assured external guard rather than implicitly by the LLM itself. Hence, we will next examine very general contracts or constraints for AI systems, and methods for enforcing them. As frameworks, we will consider ways that human behavior is constrained: that is, by ethical and legal standards, supervision, rewards and punishment, reputation, and so on. Some of this material is drawn from a previous report on controls for potentially conscious agents [210].
@@ 4.3 Ethics
Ethical and related principles are referred to as normative, and symbolic AI systems have been developed that can represent such principles and thereby perform “normative reasoning” [59]. These have been applied to trolley problems, including some that involve self-harm (e.g., throwing yourself in front of the trolley) and thereby violate the “unintended” aspect of DDE [47, 90]. It is claimed that fairly sophisticated logical treatments (e.g., intensional logics, counterfactuals, deontic modalities) are needed to represent normative scenarios, and these might be additional to what is needed for the primary functions of the system (hence, must be introduced explicitly, which will add complexity). Additionally, when normative requirements are introduced as rules in some formal notation, there is concern that they may have internal conflicts or otherwise lack wellformedness, and methods have been proposed for checking this [80]. Additional complexities arise when ethical judgments must be made against vague criteria (e.g., some medical diagnoses): is the criterion uncertain (indeterminacy [246]) or merely unknown (epistemicism [247])?
Other recent work formalizes Kant’s categorical imperative (humans must be treated as ends, not as means), which requires a treatment of causality [160], while others favor “Virtue Ethics” and “Artificial Phronesis” derived from Aristotle [236, 227, 56]. And note that we have mentioned only logical or rule-based representations for ethics, whereas game theory provides another perspective.
Application of constraints derived from general normative principles requires the guard to build a model of its world and to interpret the principles appropriately. It is unlikely that a non-AI perception system can build a model that is adequate for this task, so the model must be built by an AI perception system, or even based on that used by the LLM itself. We have seen previously that assurance for world models constructed by AI-based perception systems is challenging and currently infeasible, so only weak forms of safety assurance can be delivered by guards based on ethics and other general rules.
@@
Building an accurate and comprehensive world model is difficult in a focused application; it is next to impossible for a general-purpose tool lacking information on the possible contexts of its use.
@@
We conclude that general constraints built into guards for general-purpose AI tools cannot provide automated protection nor assurance, largely because their correct application depends on the guard building an accurate model of the world in which to interpret them, or on trusting the LLM itself to provide a suitable model (which is contrary to its generally model-free operation). It is, however, possible that, with suitable programming interfaces or suitable sensitivity to prompts, the presence of general constraints could provide useful capabilities that humans can use to add controls for specific applications built on these general-purpose tools, albeit with weak assurance, and this is a direction we would like to see pursued.
4.4 Reputation